please explain
You do not have the required permissions to view the files attached to this post.
Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.
See classesLearn live with an expert or move at your own pace. Every option includes the complete TTP study system.
Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.
with Josh Braslow, 20+ years in test prep
Self-paced EA prep. Study on your schedule.
with
Complete access from day one. Study on your schedule.
i am not happy with the way MGMAT has drafted this question... it wouldbe gr8 if stacey can help us understand it the way Mgmat wants it to be conveyed.. for me there are 2 answers...stubbornp wrote:please explain
sudhir3127 wrote:i am not happy with the way MGMAT has drafted this question... it wouldbe gr8 if stacey can help us understand it the way Mgmat wants it to be conveyed.. for me there are 2 answers...stubbornp wrote:please explain
one :
Probability of a non-nickel on first pick = (5 pennies + 4 dimes) / 15 coins = 3/5
Probability of a non-nickel on second pick = (8 non-nickel coins) / 14 coins = 4/7
thus its 3/5*4/7 = 12/35
second:
if this the Question is asking the probability of picking any combination other than NN.
Prob of NN = 6/15 * 5/14 = 1/7
Prob of not NN = 1- 1/7 = 6/7